# GALAX: Graph-Augmented Language Model for Explainable Reinforcement-Guided Subgraph Reasoning in Precision Medicine

### v1

@@ Appendix E Reinforcement Learning Generated Subgraph Details
Table 8: Model overall performances
@@
L3 $0071\scriptstyle\pm 0.0032 0.0$$013\scriptstyle\pm 0.0002 0.0$$021\scriptstyle\pm 0.0002 0.0$$011\scriptstyle\pm 0.0001 0.0$$032\scriptstyle\pm 0.0055 0.0$$0021\scriptstyle\pm 0.0037$
L3 + KG $0125\scriptstyle\pm 0.0032 0.0$$029\scriptstyle\pm 0.0003 0.0$$043\scriptstyle\pm 0.0002 0.0$$022\scriptstyle\pm 0.0001 0.0$$085\scriptstyle\pm 0.0037 0.0$$0122\scriptstyle\pm 0.0033$
L3-FT(Med) $0179\scriptstyle\pm 0.0045 0.0$$33\scriptstyle\pm 0.0064 0.011$$15\scriptstyle\pm 0.0044 0.0$$059\scriptstyle\pm 0.0023 0.0$$116\scriptstyle\pm 0.0097 0.0$$0122\scriptstyle\pm 0.0072$
L3-FT(Med) + KG $0158\scriptstyle\pm 0.0030 0.0$$058\scriptstyle\pm 0.0011 0.0$$074\scriptstyle\pm 0.0016 0.0$$038\scriptstyle\pm 0.0010 0.0$$106\scriptstyle\pm 0.0048 0.0$$0132\scriptstyle\pm 0.0040$
L3-FT(QA) $0.5250\scriptstyle\pm 0.0282$ $0.4959\scriptstyle\pm 0.0435$ $0.5094\scriptstyle\pm 0.0365$ $0.3449\scriptstyle\pm 0.0338$ $0.8889\scriptstyle\pm 0.0168$ $0.8693\scriptstyle\pm 0.0157$
L3-FT(QA) + KG $0.5185\scriptstyle\pm 0.0240$ $0.4908\scriptstyle\pm 0.0402$ $0.5038\scriptstyle\pm 0.0327$ $0.3393\scriptstyle\pm 0.0298$ $0.8794\scriptstyle\pm 0.0115$ $0.8529\scriptstyle\pm 0.0153$
G-Retriever $0.4763\scriptstyle\pm 0.0004$ $0.3929\scriptstyle\pm 0.0063$ $0.4286\scriptstyle\pm 0.0044$ $0.2757\scriptstyle\pm 0.0038$ $0.8804\scriptstyle\pm 0.0037$ $0.8550\scriptstyle\pm 0.0046$
GALAX $\mathbf{0.5472}\scriptstyle\pm\mathbf{0.0053}$ $\mathbf{0.5332}\scriptstyle\pm\mathbf{0.0031}$ $\mathbf{0.5399}\scriptstyle\pm\mathbf{0.0041}$ $\mathbf{0.3726}\scriptstyle\pm\mathbf{0.0037}$ $\mathbf{0.9249}\scriptstyle\pm\mathbf{0.0048}$ $\mathbf{0.8815}\scriptstyle\pm\mathbf{0.0033}$
Table 9: Model performances on LUAD
@@
L3 $079\scriptstyle\pm 0.0137 0.000$$5\scriptstyle\pm 0.0008 0.000$$9\scriptstyle\pm 0.0016 0.000$$05\scriptstyle\pm 0.0008 0.0$$095\scriptstyle\pm 0.0165 0.0$$0048\scriptstyle\pm 0.0082$
L3 + KG $0014\scriptstyle\pm 0.0025 0.0$$\scriptstyle\pm 0.0017 0.0011±0.00$$1\scriptstyle\pm 0.0020$ $6\scriptstyle\pm 0.0010 0.0000±0.00$$\scriptstyle\pm 0.0000 0.0$$0000\scriptstyle\pm 0.0000$
L3-FT(Med) $0091\scriptstyle\pm 0.0018 0.0$$105\scriptstyle\pm 0.0044 0.0$$079\scriptstyle\pm 0.0022 0.0$$40\scriptstyle\pm 0.0011 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(Med) + KG $0081\scriptstyle\pm 0.0071 0.0$$024\scriptstyle\pm 0.0017 0.0$$025\scriptstyle\pm 0.0002 0.0$$013\scriptstyle\pm 0.0001 0.0$$095\scriptstyle\pm 0.0165 0.0$$0048\scriptstyle\pm 0.0082$
L3-FT(QA) $0.5201\scriptstyle\pm 0.0408$ $0.4905\scriptstyle\pm 0.0532$ $0.5045\scriptstyle\pm 0.0475$ $0.3396\scriptstyle\pm 0.0433$ $0.8476\scriptstyle\pm 0.0165$ $0.8667\scriptstyle\pm 0.0218$
L3-FT(QA) + KG $0.5214\scriptstyle\pm 0.0242$ $0.4952\scriptstyle\pm 0.0432$ $0.5073\scriptstyle\pm 0.0343$ $0.3416\scriptstyle\pm 0.0314$ $0.7905\scriptstyle\pm 0.0436$ $0.8048\scriptstyle\pm 0.0541$
G-Retriever $0.4642\scriptstyle\pm 0.0181$ $0.3881\scriptstyle\pm 0.0264$ $0.4204\scriptstyle\pm 0.0233$ $0.2671\scriptstyle\pm 0.0188$ $0.8857\scriptstyle\pm 0.0000$ $0.8524\scriptstyle\pm 0.0165$
GALAX $\mathbf{0.5345}\scriptstyle\pm\mathbf{0.0185}$ $\mathbf{0.5157}\scriptstyle\pm\mathbf{0.0043}$ $\mathbf{0.5247}\scriptstyle\pm\mathbf{0.0109}$ $\mathbf{0.3581}\scriptstyle\pm\mathbf{0.0101}$ $\mathbf{0.9238}\scriptstyle\pm\mathbf{0.0436}$ $\mathbf{0.8810}\scriptstyle\pm\mathbf{0.0082}$
Table 10: Model performances on BRCA
@@
L3 $0020\scriptstyle\pm 0.0035 0.0$$017\scriptstyle\pm 0.0029 0.0$$18\scriptstyle\pm 0.0032 0.000$$9\scriptstyle\pm 0.0016 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3 + KG $0073\scriptstyle\pm 0.0068 0.0$$033\scriptstyle\pm 0.0029 0.0$$044\scriptstyle\pm 0.0038 0.0$$022\scriptstyle\pm 0.0019 0.0$$111\scriptstyle\pm 0.0192 0.0$$0056\scriptstyle\pm 0.0096$
L3-FT(Med) $0110\scriptstyle\pm 0.0086 0.0$$106\scriptstyle\pm 0.0075 0.0$$093\scriptstyle\pm 0.0068 0.0$$47\scriptstyle\pm 0.0035 0.000$$00\scriptstyle\pm 0.0000 0.0$$0111\scriptstyle\pm 0.0192$
L3-FT(Med) + KG $0149\scriptstyle\pm 0.0057 0.0$$050\scriptstyle\pm 0.0000 0.0$$071\scriptstyle\pm 0.0012 0.0$$36\scriptstyle\pm 0.0006 0.000$$00\scriptstyle\pm 0.0000 0.0$$0111\scriptstyle\pm 0.0192$
L3-FT(QA) $0.5074\scriptstyle\pm 0.0498$ $0.4856\scriptstyle\pm 0.0570$ $0.4956\scriptstyle\pm 0.0535$ $0.3336\scriptstyle\pm 0.0491$ $\mathbf{0.8889}\scriptstyle\pm\mathbf{0.0509}$ $0.8389\scriptstyle\pm 0.0096$
L3-FT(QA) + KG $0.4856\scriptstyle\pm 0.0395$ $0.4656\scriptstyle\pm 0.0436$ $0.4751\scriptstyle\pm 0.0412$ $0.3142\scriptstyle\pm 0.0354$ $0.8778\scriptstyle\pm 0.0192$ $0.8222\scriptstyle\pm 0.0347$
G-Retriever $0.4414\scriptstyle\pm 0.0099$ $0.3772\scriptstyle\pm 0.0010$ $0.4062\scriptstyle\pm 0.0034$ $0.2607\scriptstyle\pm 0.0032$ $0.8667\scriptstyle\pm 0.0000$ $\mathbf{0.8667}\scriptstyle\pm\mathbf{0.0000}$
@@
Table 11: Model performances on COAD/READ

### v2

@@ D.3 Baseline Models
Other Models
RoG, SubgraphRAG, and GNN-RAG are evidence-subgraph-dependent methods that share a common assumption: a small, clean, and reliable subgraph exists from which a retriever module can be trained to provide reasoning paths to the LLM. However, this assumption breaks down in large-scale biomedical knowledge graphs, which are inherently noisier and lack validated reasoning paths. In the Target-QA setting, each question is paired with a one-hop disease–protein subgraph and a one-hop PPI neighborhood extracted from BioMedGraphica. These subgraphs are extremely noisy and large (containing thousands of nodes and edges), and lack ground-truth mechanistic subgraphs, making supervised retrieval difficult. Specifically, RoG, SubgraphRAG, and GNN-RAG require ground-truth reasoning paths for training. Since such annotations are unavailable in our setting, we employ ChatGPT 5.1 to generate candidate reasoning paths relevant to the annotated cell lines, incorporating disease context and multi-omic profiles. However, this GPT-derived supervision is inherently noisy and biologically incomplete, limiting retriever effectiveness. All baselines use the fine-tuned Llama3-8B-Instruct as the backbone LLM.
@@ D.4 Results
D.5 Generalization to Unseen Cancer Types
To assess how well GALAX generalizes across cancer types, we constructed three holdout sets in which all cell lines under same TCGA codes were removed during training and used only for evaluation (see Table 12). This design forces the model to generate CRISPR target predictions for cancer types not observed during optimization, providing a direct measure of cross-cancer transfer and robustness to unseen molecular contexts.
As expected, performance decreases when entire cancer types are held out, since no fine-tuning data from those cancers is available. Nevertheless, the results in Table 13 show that GALAX maintains stable performance across the three held-out cancer groups. Precision, recall, F1, and Jaccard values remain consistent, with an average F1 of $0.4862$ and Hit@10 above $0.85$. While these scores are lower than the full-data GALAX model, the gap is moderate, indicating that the model retains the ability to infer relevant targets even when the TCGA code of the input cell line has never appeared during training, as long as it has been well pretrained with sufficient samples. This demonstrates that GALAX does not overfit to specific cancer labels and can transfer its reasoning across diverse cell-line backgrounds, addressing concerns regarding generalization to new cancer types.
Table 12: Held-out cancer type fold assignments
Holdout Set Train Test Train/Test TCGA Codes Hold-out Cancer TCGA Codes
Holdout Set 1 238 (65.6%) 125 (34.4%) 20 / 10 COAD/READ, DLBC, KIRC, LCML, LUSC
PAAD, PRAD, SARC, SKCM, STAD
Holdout Set 2 289 (79.6%) 74 (20.4%) 21 / 9 BLCA, GBM, LAML, LGG, LIHC
MESO, MM, NB, UCEC
Holdout Set 3 220 (60.6%) 143 (39.4%) 21 / 9 ALL, BRCA, ESCA, HNSC, LUAD
MB, OV, SCLC, THCA
Table 13: GALAX performance on unseen cancer types under held-out evaluation
Holdout Set Precision $\uparrow$ Recall $\uparrow$ F1 $\uparrow$ Jaccard $\uparrow$ Hit@5 $\uparrow$ Hit@10 $\uparrow$
Holdout Set 1 $0.4931$ $0.4501$ $0.4699$ $0.3100$ $0.9024$ $0.8688$
Holdout Set 2 $0.5124$ $0.5151$ $0.5123$ $0.3465$ $0.8703$ $0.8581$
Holdout Set 3 $0.4832$ $0.4715$ $0.4765$ $0.3161$ $0.8993$ $0.8448$
Average $0.4962$ $0.4789$ $0.4862$ $0.3242$ $0.8907$ $0.8572$
GALAX $0.5472\scriptstyle\pm 0.0053$ $0.5332\scriptstyle\pm 0.0031$ $0.5399\scriptstyle\pm 0.0041$ $0.3726\scriptstyle\pm 0.0037$ $0.9249\scriptstyle\pm 0.0048$ $0.8815\scriptstyle\pm 0.0033$
D.6 External datasets
To examine generalization beyond Target-QA, we performed an external evaluation on the pediatric cancer dataset from PedDep [^6] https://peddep.org/, which contains CRISPR-based gene effect profiles and cell-line annotations for 31 pediatric cancer samples. The data were processed following the same steps as Target-QA. Due to the limited sample size, we did not train or finetune on this dataset and instead evaluated all models in a zero-shot setting. Given that all samples are cancerous, the M2T baseline cannot be applied. As shown in Table 14, baseline graph models and language-model variants still have near-zero performance. Graph-augmented methods such as RoG, SubgraphRAG, and GNN-RAG yield moderate gains, but GALAX attains the highest precision, recall, F1, and Jaccard scores, along with notable improvements in Hit@5 and Hit@10 finetuned on Target-QA with multi-omics and disease-related proteins information. These results demonstrate that GALAX can extend its target-prediction ability to external cancer datasets without retraining, providing strong evidence of its robustness and transferability beyond the Target-QA benchmark.
Table 14: Model performance on PedDep cancer dataset
GAT $005\scriptstyle\pm 0.0008 0.000$$5\scriptstyle\pm 0.0008 0.000$$5\scriptstyle\pm 0.0008 0.000$$2\scriptstyle\pm 0.0004 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(Med) + Omics $0144\scriptstyle\pm 0.0081 0.0$$114\scriptstyle\pm 0.0049 0.0$$099\scriptstyle\pm 0.0028 0.0$$050\scriptstyle\pm 0.0014 0.0$$210\scriptstyle\pm 0.0421 0.0$$0109\scriptstyle\pm 0.0218$
L3-FT(Med) + Omics + KG $0167\scriptstyle\pm 0.0161 0.0$$060\scriptstyle\pm 0.0047 0.0$$068\scriptstyle\pm 0.0049 0.0$$035\scriptstyle\pm 0.0025 0.0$$072\scriptstyle\pm 0.0143 0.0$$0073\scriptstyle\pm 0.0145$
L3 + Omics $0054\scriptstyle\pm 0.0108 0.0$$011\scriptstyle\pm 0.0021 0.0$$18\scriptstyle\pm 0.0036 0.000$$9\scriptstyle\pm 0.0018 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3 + Omics + KG $0192\scriptstyle\pm 0.0188 0.0$$032\scriptstyle\pm 0.0022 0.0$$050\scriptstyle\pm 0.0035 0.0$$026\scriptstyle\pm 0.0018 0.0$$142\scriptstyle\pm 0.0165 0.0$$0214\scriptstyle\pm 0.0247$
L3-FT(QA) + Omics $0.2608\scriptstyle\pm 0.0268$ $5\scriptstyle\pm 0.0252 0.2606±0.02$$6\scriptstyle\pm 0.0260$ $0.1517\scriptstyle\pm 0.0152$ $0.5810\scriptstyle\pm 0.0330$ $0.4524\scriptstyle\pm 0.0502$
L3-FT(QA) + Omics + KG $0.2619\scriptstyle\pm 0.0075$ $0.2552\scriptstyle\pm 0.0095$ $0.2584\scriptstyle\pm 0.0086$ $0.1489\scriptstyle\pm 0.0055$ $0.5048\scriptstyle\pm 0.0165$ $0.4952\scriptstyle\pm 0.0218$
G-Retriever + pre-GAT $0.2624\scriptstyle\pm 0.0212$ $0.2610\scriptstyle\pm 0.0206$ $0.2617\scriptstyle\pm 0.0209$ $0.1522\scriptstyle\pm 0.0121$ $0.5143\scriptstyle\pm 0.0495$ $0.4524\scriptstyle\pm 0.0825$
RoG $0.2730\scriptstyle\pm 0.0050$ $0.2667\scriptstyle\pm 0.0092$ $0.2697\scriptstyle\pm 0.0071$ $0.1566\scriptstyle\pm 0.0047$ $0.5143\scriptstyle\pm 0.0286$ $0.4714\scriptstyle\pm 0.0623$
SubgraphRAG $0.2736\scriptstyle\pm 0.0091$ $0.2690\scriptstyle\pm 0.0095$ $0.2712\scriptstyle\pm 0.0092$ $0.1579\scriptstyle\pm 0.0062$ $0.5619\scriptstyle\pm 0.0165$ $0.5286\scriptstyle\pm 0.0571$
GNN-RAG $0.2760\scriptstyle\pm 0.0092$ $0.2700\scriptstyle\pm 0.0094$ $0.2728\scriptstyle\pm 0.0093$ $0.1589\scriptstyle\pm 0.0064$ $0.5714\scriptstyle\pm 0.0495$ $0.5333\scriptstyle\pm 0.0360$
GALAX $0.2914\scriptstyle\pm 0.0115$ $0.2889\scriptstyle\pm 0.0131$ $0.2901\scriptstyle\pm 0.0123$ $0.1703\scriptstyle\pm 0.0086$ $0.6357\scriptstyle\pm 0.0589$ $0.5179\scriptstyle\pm 0.0513$
GALAX (Qwen2.5-7B) $\mathbf{0.2921}\scriptstyle\pm\mathbf{0.0055}$ $\mathbf{0.2895}\scriptstyle\pm\mathbf{0.0050}$ $\mathbf{0.2908}\scriptstyle\pm\mathbf{0.0053}$ $\mathbf{0.1708}\scriptstyle\pm\mathbf{0.0038}$ $\mathbf{0.6667}\scriptstyle\pm\mathbf{0.0595}$ $\mathbf{0.5381}\scriptstyle\pm\mathbf{0.0705}$
@@ E.1 Enrichment analysis
E.2 Human evaludation and LLM as judge
To further assess the biological relevance of the generated subgraphs, we conducted a comprehensive evaluation involving both human domain experts and Large Language Models (LLMs). The human evaluation panel consisted of three bioinformaticians, denoted as $\mathbf{h}_{1},\mathbf{h}_{2},$ and $\mathbf{h}_{3}$. In parallel, we used two advanced LLMs as automated evaluators ($\mathbf{m}_{1},\mathbf{m}_{2}$): ChatGPT-5.1 and Gemini-3.0 Pro. Both the human experts and the LLMs were provided with the generated subgraphs, corresponding gene and pathway enrichment analysis results to facilitate the assessment of biological plausibility.
Table 15: Human Evaluation Scores
Sample ID TCGA Code $\mathbf{h}_{1}$ $\mathbf{h}_{2}$ $\mathbf{h}_{3}$ $\mathbf{m}_{1}$ $\mathbf{m}_{2}$ Mean $\pm$ Std
ACH-000860 LUAD 4 5 5 5 5 4.80 $\pm$ 0.45
ACH-000054 SRCA 3 5 3 4 4 3.80 $\pm$ 0.84
ACH-000001 OV 3 3 2 4 2 2.80 $\pm$ 0.84
ACH-000219 SKCM 4 5 2 4 3 3.60 $\pm$ 1.14
ACH-000070 ALL 5 5 5 4 5 4.80 $\pm$ 0.45
ACH-000092 MESO 3 3 2 3 3 2.80 $\pm$ 0.45
ACH-000817 $M$ 5 5 5 3 3 4.20 $\pm$ 1.10
ACH-000649 KIRC 4 4 5 5 5 4.60 $\pm$ 0.55
ACH-000018 BLCA 3 3 3 3 3 3.00 $\pm$ 0.00
ACH-000864 UCEC 3 4 2 4 3 3.20 $\pm$ 0.84
Overall - 3.7 4.2 3.4 3.9 3.6 $\mathbf{3.76}\pm 0.80$
Each expert independently reviewed ten subgraph examples shown in Table 15 and assigned a score from 1 to 5 based on the degree of correspondence between the subgraph and the known biology of the associated TCGA cancer type. The scoring rubric was defined as follows:
  • 5: Highly related — strong and clear match to hallmark pathways and well-established features of the cancer type.
  • 4: Related — clear relationship to the cancer type but less comprehensive or slightly mixed.
  • 3: Moderately related — some relevant elements present, but mixed with non-specific findings.
  • 2: Not related — only weak or indirect connection to the cancer type.
  • 1: Not related at all — no meaningful alignment with the biology of the cancer type.
As shown in Table 15, several samples were evaluated as related or highly related to their TCGA cancer types, indicating that the subgraphs generated by the model capture meaningful molecular features. The LUAD sample ACH-000860 received one of the highest average scores, as its subgraph highlighted pathways such as regulation of phosphatidylinositol 3-kinase signaling and positive regulation of kinase activity, involving genes like EGFR, PTK2, and EPHB4 that are central to lung adenocarcinoma biology. The ALL sample ACH-000070 was also rated highly due to the presence of TCF3 and PBX1, whose fusion is a well-established driver event in Acute Lymphoblastic Leukemia.
Table 16: Model Overall performance on Target-QA
@@ E.2 Human evaludation and LLM as judge
L3 + Omics $0071\scriptstyle\pm 0.0032 0.0$$013\scriptstyle\pm 0.0002 0.0$$021\scriptstyle\pm 0.0002 0.0$$011\scriptstyle\pm 0.0001 0.0$$032\scriptstyle\pm 0.0055 0.0$$0021\scriptstyle\pm 0.0037$
L3 + Omics + KG $0125\scriptstyle\pm 0.0032 0.0$$029\scriptstyle\pm 0.0003 0.0$$043\scriptstyle\pm 0.0002 0.0$$022\scriptstyle\pm 0.0001 0.0$$085\scriptstyle\pm 0.0037 0.0$$0122\scriptstyle\pm 0.0033$
L3-FT(Med) + Omics $0179\scriptstyle\pm 0.0045 0.0$$33\scriptstyle\pm 0.0064 0.011$$15\scriptstyle\pm 0.0044 0.0$$059\scriptstyle\pm 0.0023 0.0$$116\scriptstyle\pm 0.0097 0.0$$0122\scriptstyle\pm 0.0072$
L3-FT(Med) + Omics + KG $0158\scriptstyle\pm 0.0030 0.0$$058\scriptstyle\pm 0.0011 0.0$$074\scriptstyle\pm 0.0016 0.0$$038\scriptstyle\pm 0.0010 0.0$$106\scriptstyle\pm 0.0048 0.0$$0132\scriptstyle\pm 0.0040$
L3-FT(QA) + Omics $0.5250\scriptstyle\pm 0.0282$ $0.4959\scriptstyle\pm 0.0435$ $0.5094\scriptstyle\pm 0.0365$ $0.3449\scriptstyle\pm 0.0338$ $0.8889\scriptstyle\pm 0.0168$ $0.8693\scriptstyle\pm 0.0157$
L3-FT(QA) + Omics + KG $0.5185\scriptstyle\pm 0.0240$ $0.4908\scriptstyle\pm 0.0402$ $0.5038\scriptstyle\pm 0.0327$ $0.3393\scriptstyle\pm 0.0298$ $0.8794\scriptstyle\pm 0.0114$ $0.8529\scriptstyle\pm 0.0153$
G-Retriever + pre-GAT $0.4763\scriptstyle\pm 0.0004$ $0.3929\scriptstyle\pm 0.0063$ $0.4286\scriptstyle\pm 0.0044$ $0.2757\scriptstyle\pm 0.0038$ $0.8804\scriptstyle\pm 0.0037$ $0.8550\scriptstyle\pm 0.0046$
RoG $0.5248\scriptstyle\pm 0.0134$ $0.4726\scriptstyle\pm 0.0445$ $0.4924\scriptstyle\pm 0.0323$ $0.3338\scriptstyle\pm 0.0267$ $0.8593\scriptstyle\pm 0.0318$ $0.8450\scriptstyle\pm 0.0350$
SubgraphRAG $0.5280\scriptstyle\pm 0.0044$ $0.4617\scriptstyle\pm 0.0027$ $0.4860\scriptstyle\pm 0.0033$ $0.3269\scriptstyle\pm 0.0024$ $0.8624\scriptstyle\pm 0.0120$ $0.8476\scriptstyle\pm 0.0167$
GNN-RAG $0.5258\scriptstyle\pm 0.0126$ $0.4735\scriptstyle\pm 0.0190$ $0.4935\scriptstyle\pm 0.0168$ $0.3345\scriptstyle\pm 0.0134$ $0.8656\scriptstyle\pm 0.0302$ $0.8323\scriptstyle\pm 0.0205$
GALAX $\mathbf{0.5472}\scriptstyle\pm\mathbf{0.0053}$ $0.5332\scriptstyle\pm 0.0031$ $0.5399\scriptstyle\pm 0.0041$ $0.3726\scriptstyle\pm 0.0037$ $\mathbf{0.9249}\scriptstyle\pm\mathbf{0.0048}$ $0.8815\scriptstyle\pm 0.0033$
GALAX (Qwen2.5-7B) $0.5445\scriptstyle\pm 0.0114$ $\mathbf{0.5405}\scriptstyle\pm\mathbf{0.0101}$ $\mathbf{0.5422}\scriptstyle\pm\mathbf{0.0104}$ $\mathbf{0.3744}\scriptstyle\pm\mathbf{0.0098}$ $0.9079\scriptstyle\pm 0.0084$ $\mathbf{0.8841}\scriptstyle\pm\mathbf{0.0126}$
Table 17: Model performance on LUAD
@@
L3 + Omics $079\scriptstyle\pm 0.0137 0.000$$5\scriptstyle\pm 0.0008 0.000$$9\scriptstyle\pm 0.0016 0.000$$05\scriptstyle\pm 0.0008 0.0$$095\scriptstyle\pm 0.0165 0.0$$0048\scriptstyle\pm 0.0082$
L3 + Omics + KG $0014\scriptstyle\pm 0.0025 0.0$$\scriptstyle\pm 0.0016 0.0011±0.00$$1\scriptstyle\pm 0.0020$ $6\scriptstyle\pm 0.0010 0.0000±0.00$$\scriptstyle\pm 0.0000 0.0$$0000\scriptstyle\pm 0.0000$
L3-FT(Med) + Omics $0091\scriptstyle\pm 0.0018 0.0$$105\scriptstyle\pm 0.0044 0.0$$079\scriptstyle\pm 0.0022 0.0$$40\scriptstyle\pm 0.0011 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3-FT(Med) + Omics + KG $0081\scriptstyle\pm 0.0071 0.0$$024\scriptstyle\pm 0.0016 0.0$$025\scriptstyle\pm 0.0002 0.0$$013\scriptstyle\pm 0.0001 0.0$$095\scriptstyle\pm 0.0165 0.0$$0048\scriptstyle\pm 0.0082$
L3-FT(CRISPR) + Omics $0.5201\scriptstyle\pm 0.0408$ $0.4905\scriptstyle\pm 0.0532$ $0.5045\scriptstyle\pm 0.0475$ $0.3396\scriptstyle\pm 0.0433$ $0.8476\scriptstyle\pm 0.0165$ $0.8667\scriptstyle\pm 0.0218$
L3-FT(CRISPR) + Omics + KG $0.5214\scriptstyle\pm 0.0242$ $0.4952\scriptstyle\pm 0.0432$ $0.5073\scriptstyle\pm 0.0343$ $0.3416\scriptstyle\pm 0.0314$ $0.7905\scriptstyle\pm 0.0436$ $0.8048\scriptstyle\pm 0.0541$
G-Retriever + pre-GAT $0.4642\scriptstyle\pm 0.0181$ $0.3881\scriptstyle\pm 0.0264$ $0.4204\scriptstyle\pm 0.0233$ $0.2671\scriptstyle\pm 0.0188$ $0.8857\scriptstyle\pm 0.0000$ $0.8524\scriptstyle\pm 0.0165$
RoG $0.5213\scriptstyle\pm 0.0227$ $0.4562\scriptstyle\pm 0.0848$ $0.4793\scriptstyle\pm 0.0630$ $0.3228\scriptstyle\pm 0.0530$ $0.8095\scriptstyle\pm 0.0436$ $0.8238\scriptstyle\pm 0.0218$
SubgraphRAG $0.5123\scriptstyle\pm 0.0105$ $0.4448\scriptstyle\pm 0.0386$ $0.4684\scriptstyle\pm 0.0279$ $0.3114\scriptstyle\pm 0.0237$ $0.8190\scriptstyle\pm 0.0165$ $0.8238\scriptstyle\pm 0.0082$
GNN-RAG $334\scriptstyle\pm 0.0225 0.505$$052\scriptstyle\pm 0.0170$ $0.5165\scriptstyle\pm 0.0161$ $0.3563\scriptstyle\pm 0.0127$ $0.7905\scriptstyle\pm 0.0719$ $0.7571\scriptstyle\pm 0.0623$
GALAX $0.5345\scriptstyle\pm 0.0185$ $0.5157\scriptstyle\pm 0.0043$ $0.5247\scriptstyle\pm 0.0109$ $0.3581\scriptstyle\pm 0.0101$ $\mathbf{0.9238}\scriptstyle\pm\mathbf{0.0436}$ $\mathbf{0.8810}\scriptstyle\pm\mathbf{0.0082}$
GALAX (Qwen2.5-7B) $\mathbf{0.5475}\scriptstyle\pm\mathbf{0.0019}$ $\mathbf{0.5462}\scriptstyle\pm\mathbf{0.0111}$ $\mathbf{0.5465}\scriptstyle\pm\mathbf{0.0050}$ $\mathbf{0.3778}\scriptstyle\pm\mathbf{0.0046}$ $0.9048\scriptstyle\pm 0.0165$ $0.8667\scriptstyle\pm 0.0082$
Table 18: Model performance on BRCA
@@
L3 + Omics $0020\scriptstyle\pm 0.0035 0.0$$017\scriptstyle\pm 0.0029 0.0$$18\scriptstyle\pm 0.0032 0.000$$9\scriptstyle\pm 0.0016 0.000$$\scriptstyle\pm 0.0000 0.0000±0$$00\scriptstyle\pm 0.0000$
L3 + Omics + KG $0073\scriptstyle\pm 0.0068 0.0$$033\scriptstyle\pm 0.0029 0.0$$044\scriptstyle\pm 0.0038 0.0$$022\scriptstyle\pm 0.0019 0.0$$111\scriptstyle\pm 0.0192 0.0$$0056\scriptstyle\pm 0.0096$
L3-FT(Med) + Omics $0110\scriptstyle\pm 0.0086 0.0$$106\scriptstyle\pm 0.0075 0.0$$093\scriptstyle\pm 0.0068 0.0$$47\scriptstyle\pm 0.0035 0.000$$00\scriptstyle\pm 0.0000 0.0$$0111\scriptstyle\pm 0.0192$
L3-FT(Med) + Omics + KG $0149\scriptstyle\pm 0.0057 0.0$$050\scriptstyle\pm 0.0000 0.0$$071\scriptstyle\pm 0.0012 0.0$$36\scriptstyle\pm 0.0006 0.000$$00\scriptstyle\pm 0.0000 0.0$$0111\scriptstyle\pm 0.0192$
L3-FT(CRISPR) + Omics $0.5074\scriptstyle\pm 0.0498$ $0.4856\scriptstyle\pm 0.0570$ $0.4956\scriptstyle\pm 0.0535$ $0.3336\scriptstyle\pm 0.0491$ $\mathbf{0.8889}\scriptstyle\pm\mathbf{0.0509}$ $0.8389\scriptstyle\pm 0.0096$
L3-FT(CRISPR) + Omics + KG $0.4856\scriptstyle\pm 0.0395$ $0.4656\scriptstyle\pm 0.0436$ $0.4751\scriptstyle\pm 0.0412$ $0.3142\scriptstyle\pm 0.0354$ $0.8778\scriptstyle\pm 0.0192$ $0.8222\scriptstyle\pm 0.0347$
G-Retriever + pre-GAT $0.4414\scriptstyle\pm 0.0099$ $0.3772\scriptstyle\pm 0.0010$ $0.4062\scriptstyle\pm 0.0034$ $0.2607\scriptstyle\pm 0.0032$ $0.8667\scriptstyle\pm 0.0000$ $\mathbf{0.8667}\scriptstyle\pm\mathbf{0.0000}$
RoG $0.4791\scriptstyle\pm 0.0575$ $0.4311\scriptstyle\pm 0.0721$ $0.4489\scriptstyle\pm 0.0658$ $0.2999\scriptstyle\pm 0.0528$ $0.7667\scriptstyle\pm 0.0577$ $0.7611\scriptstyle\pm 0.1110$
SubgraphRAG $0.4708\scriptstyle\pm 0.0317$ $0.3917\scriptstyle\pm 0.0376$ $0.4196\scriptstyle\pm 0.0339$ $0.2742\scriptstyle\pm 0.0271$ $0.7556\scriptstyle\pm 0.0839$ $0.7333\scriptstyle\pm 0.1014$
GNN-RAG $0.4787\scriptstyle\pm 0.0453$ $0.4389\scriptstyle\pm 0.0584$ $0.4543\scriptstyle\pm 0.0531$ $0.3025\scriptstyle\pm 0.0428$ $0.8444\scriptstyle\pm 0.0385$ $0.8222\scriptstyle\pm 0.0674$
@@
GALAX (Qwen2.5-7B) $0.5171\scriptstyle\pm 0.0474$ $0.5206\scriptstyle\pm 0.0419$ $0.5188\scriptstyle\pm 0.0448$ $0.3532\scriptstyle\pm 0.0392$ $0.8556\scriptstyle\pm 0.0385$ $0.8000\scriptstyle\pm 0.0764$
Table 19: Model performance on COAD/READ